### v1

@@ 1 Introduction
To make progress, we recall the definition of the ensemble. Take a product of correlation functions and take the average. This is called a multi-copy observable [7]. In an ensemble, the average of a product does not factorize into the product of the averages. Computing this is complicated in general, as this object depends on the OPE coefficients and the spectrum, both being random variables. Moreover, the correlation function depends on the conformal weights in a non-algebraic manner. A tensor model has been designed to deal with this general situation [9, 13]. However, our goal is less ambitious as we work at large $c$ where things simplify significantly. For one thing, the wormhole that signals the random matrix nature of the spectrum is order one in $c$ [1].
@@
Computing multi-copy observables therefore reduces to a simpler problem. As the correlation functions depend on the OPE coefficients algebraically, conformal blocks can essentially be moved outside of the average. One therefore only needs to know about quantities that are of the schematic form $\overline{C_{ijk}^{n}}$, where $C_{ijk}$ is the OPE coefficient and the overline denotes averaging. These are multivariate statistical moments for the OPE coefficients, or OPE moments for short.
@@
While defining an ensemble from an intrinsically CFT point of view has been attempted [9, 21], establishing the agreement with 3D gravity is still an effort in progress. [^1] Certain OPE moments are in some sense universal and obtainable from bootstrap [22, 6], and the corresponding gravity answers are known to agree with the bootstrap results [7, 5]. In these situations, the agreement can be understood by uplifting the bootstrap procedure to 3D, as explained in [18], Section 8.4. We circumvent this issue by defining the ensemble via 3D gravity. To do this, we essentially follow [7] and [8, 11]. The OPE moment is defined by finding wormholes connecting asymptotic boundaries with the topology of a three-punctured sphere. Each three-punctured sphere represents an OPE coefficient. In the bulk, Virasoro Wilson lines connect the asymptotic punctures with the same labels. At large $c$, only certain manifolds with Wilson lines inserted in certain ways have on-shell configurations, and we call them hyperbolic manifolds. We should emphasize that the Wilson line configuration is part of the definition of hyperbolicity, as Virasoro Wilson lines encode topological information (unlike matter Wilson lines). Geometrically, they become spacetime bananas [23, 24] at large $c$ if the states connected by the Wilson lines are above the black hole threshold, or conical defects if they are below the threshold [25, 26]. In this work, we do not include any states below the threshold in the ensemble.
@@
Now that we have a definition of the OPE moments, we can go back to the multi-copy observables and pose the question of crossing symmetry: do we get the same answer for a multi-copy observable regardless of the choice of the OPE channel for each copy? From the perspective of constructing the ensemble, this is a strong requirement. For example, it forces the ensemble to be non-Gaussian [5]. From the gravity perspective, however, this seems intuitive: diffeomorphism of gravity ensures that no channel is preferred. The goal of the current work is to establish this explicitly and to clarify several points that will help manifest the crossing symmetry.
@@ 5 Discussion
We have focused on closed CFTs. The connection between 3D gravity and CFT ensembles generalizes to BCFTs [18, 19], where the ensemble gains new random variables and the bulk gains new dynamical objects. At finite $c$, the bulk theory is an open-closed extension of Virasoro TQFT [19, 20]. A BCFT needs to satisfy six independent bootstrap conditions (crossing symmetries) [36, 37], so it is natural to ask whether the main results of this work generalize to the BCFT ensemble. For example, this will require a generalization of the notation of pseudo-hyperbolicity to manifolds with end-of-the-world branes. It would be interesting to do this explicitly.

### v2

@@ 1 Introduction
To make progress, we recall the definition of the ensemble. Take a product of correlation functions and take the average. This is called a multi-copy observable [7]. In an ensemble, the average of a product does not generally factorize into the product of the averages. Computing this is complicated in general, as this object depends on the OPE coefficients and the spectrum, both being random variables. Moreover, the correlation function depends on the conformal weights in a non-algebraic manner. A tensor model has been designed to deal with this general situation [9, 13].
@@
However, our goal is less ambitious as we work at large $c$ where things simplify significantly. For one thing, the wormhole that signals the random matrix nature of the spectrum is order one in $c$ [1], allowing us to treat the spectrum as fixed rather than random. Computing multi-copy observables then reduces to a more manageable problem. As the correlation functions depend on the OPE coefficients algebraically, conformal blocks can essentially be moved outside of the average. One therefore only needs to know about quantities that are of the schematic form $\overline{C_{ijk}^{n}}$, where $C_{ijk}$ is the OPE coefficient and the overline denotes averaging. These are multivariate statistical moments for the OPE coefficients, or OPE moments for short.
@@
While defining an ensemble from an intrinsically CFT point of view has been attempted [9, 21], establishing the agreement with 3D gravity is still an effort in progress. [^1] Certain OPE moments are in some sense universal and obtainable from bootstrap [22, 6], and the corresponding gravity answers are known to agree with the bootstrap results [7, 5]. In these situations, the agreement can be understood by uplifting the bootstrap procedure to 3D, as explained in [18], Section 8.4. We circumvent this issue by defining the ensemble via 3D gravity. To do this, we follow [8, 11]. The OPE moment is defined by finding wormholes connecting asymptotic boundaries with the topology of a three-punctured sphere. Each three-punctured sphere represents an OPE coefficient. In the bulk, Virasoro Wilson lines connect the asymptotic punctures with the same labels. In the large-$c$ limit, only certain manifolds with Wilson lines inserted in certain ways contribute. These are hyperbolic manifolds, defined to be the set of manifolds with on-shell configurations. We should emphasize that the Wilson line configuration is part of the definition of hyperbolicity, as Virasoro Wilson lines encode topological information (unlike matter Wilson lines). Geometrically, they become spacetime bananas [23, 24] at large $c$ if the states connected by the Wilson lines are above the black hole threshold, or conical defects if they are below the threshold [25, 26]. In this work, we do not include any states below the threshold in the ensemble.
@@
Now that we have a definition of the OPE moments, we can go back to the multi-copy observables and pose the question of crossing symmetry: do we get the same answer for a multi-copy observable regardless of the choice of the OPE channel for each copy? From the perspective of constructing the ensemble, this is a strong requirement. For example, it forces the ensemble to be non-Gaussian [5]. From the gravity perspective, however, this seems intuitive: diffeomorphism of gravity should ensure that no channel is preferred. The goal of the current work is to establish this explicitly and to clarify several points that will help manifest the crossing symmetry.
@@
Note added.
The ideas of “ensemble crossing symmetry” and “joining” were first presented publicly in multiple talks by the authors of [27] and were also communicated to me privately prior to the submission of version 1 of the present preprint.
@@ 5 Discussion
Along a different line, the idea of using crossing symmetry to define an ensemble has been employed in the tensor model of [9]. By construction, the model is manifestly crossing symmetric. However, it is still a conjecture that the model produces each 3D manifold exactly once [13]. The statement that crossing symmetry holds if each 3D manifold is counted with weight one, which we demonstrate in this work, is therefore a piece of supporting evidence for the conjecture.
We have focused on closed CFTs. The connection between 3D gravity and CFT ensembles generalizes to BCFTs [18, 19], where the ensemble gains new random variables and the bulk gains new dynamical objects. At finite $c$, the bulk theory is an open-closed extension of Virasoro TQFT [19, 20]. A BCFT needs to satisfy six independent bootstrap conditions (crossing symmetries) [37, 38], so it is natural to ask whether the main results of this work generalize to the BCFT ensemble. For example, this will require a generalization of the notation of pseudo-hyperbolicity to manifolds with end-of-the-world branes. It would be interesting to establish this explicitly.